Algebra

If b

Rayon
7 Years agoGrade

4 Answers

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
Please provide full question as the question is posted here is incomplete.
You can also attach an image
Rayon
7 Years ago
If b is less than 0 and x1 and x2 are roots of equation 2x² + 6x + b = 0. And x1/x2 + x2/x1 is less than K. Find value is k
Deepak Kumar Shringi
7 Years ago
incomplete question
Arun
7 Years ago
Dear Rayon
 
x1 /x2 + x2/ x1 = (x1+ x2)^2 – 2 x1* x2 /x1 * x2 = 18/b – 2
 
Since b is less than 0 first term will be negativeSo maximum value of above will be -2